Q1) Let S ≡ 4x2 −9y2−36 = 0 and S' ≡ y2 −4x = 0 be two conics. Under what

conditions on k, will the conic S+kS' = 0 represent:

i) an ellipse?

ii) a hyperbola?


Q2) Find the section of the conicoid x2/2-y2/3= 2z by the plane x−2y+z = 1. What

conic does this section represent? Justify your answer. 


Q3)If 1,1/2,0 are direction ratios of a line, then the line makes an angle of 90◦ with the

x-axis, an angle of 60◦ with the y-axis, and is parallel to the z-axis.


Q4)If a cone has three mutually perpendicular generators then its reciprocal cone has

three mutually perpendicular tangent planes.


1
Expert's answer
2021-05-17T16:02:49-0400

1)

The conic "Ax^2+Bxy+Cy^2+Dx+Ey+F=0"

is an ellipse if "B^2<4AC" and a hyperbola if "B^2-4AC>0"

"S+kS'=4x^2+0xy+(k-9)y^2-4kx+0y-36=0"

is an ellipse if

02<4(4)(k-9)

0<16k-144

144<16k

k>9

an a hyperbola if

02-4(4)(k-9)>0

0-16k+144>0

-16k>-144

k<9



2)

"\frac{x^2}{2}-\frac{y^2}{3}=2z \\
x-2y+z=1\\
z=1-x+2y
\\
\frac{x^2}{2}-\frac{y^2}{3}=2(1-x+2z)\\
(\sqrt{3}x+2\sqrt{3})^2-(\sqrt{2}y+3\sqrt{2})^2=42\\
\frac{(x+2)^2}{14}-\frac{(y+3)^2}{21}=1"


its a hyperbola


3)

If 1,1/2,0 are direction ratios of a line, then the line makes an angle of 90◦ with the

x-axis, an angle of 60◦ with the y-axis, and is parallel to the z-axis.

=True


4)

(l,m,n) are DC's of axes then

"l=m=n=\cos\theta=\left(\frac13\right)^\frac12"


"\tanθ=(2)^\frac12\implies θ=\arctan(2^\frac12)"